OpenStudy (anonymous):

Using the properties of exponents and radicals, design at least three different equivalent forms of x². You must show how each one can be simplified back to x² in two or more steps. Stretch your mind and get creative! Keep in mind that something too simple, like x • x would not be acceptable since it takes only one step to convert it to x².

OpenStudy (anonymous):

sqrt(x^4) ?

OpenStudy (anonymous):

I don't understand what it's asking me to do. Like I need 3 different equations that can be simplified down to x^2 but in 2+ steps

OpenStudy (owlcoffee):

It's like building a exponential equation exercise, in other words, let's begin with x^2 : \[x ^{2}\] Nice... Now, I think I'll add some division (wich implied the sustraction of exponents) \[\frac{ x ^{4} }{ x ^{2} }\] Looks fun... I'll add some radicals. \[\frac{ \sqrt[]{x ^{8}} }{ \sqrt{x ^{4}} }\] I'll create another one, a multiplication that makes everything 1. meaning i'm searching for a exponent equal to zero: \[(\frac{ \sqrt{x^{8}} }{ \sqrt{x ^{4}} })(\frac{ y ^{3} }{ y ^{3} })\] To make thing more fun, I'll transform each of y^3 in radicals and get rid of the parenthesis: \[\frac{ \sqrt[3]{y ^{6}}\sqrt{x ^{8}} }{ \sqrt[4]{y ^{12}}\sqrt{x ^{4}} }\] And that's just an example, just play around with the properties.

OpenStudy (anonymous):

Thank you so much